How many moles of potassium dihydrogen phosphate (KH2PO4) are in
67.9
67.9
g of the compound
1
Expert's answer
2019-07-22T07:22:47-0400

m (KH2PO4)=67, 9 g

n(KH2PO4)-?

n(KH2PO4)= m(KH2PO4)/ M(KH2PO4)=67,9/136.086=0.499 mol≈0,5 mol